OpenStudy (anonymous):

Solve for x: 4 − (x + 2) < −3(x + 4) x > −9 x < −7 x > −7 x < −9 I keep getting x < -9 but then I got it wrong. I am aloud to retake this quiz but I need help getting the answer.

OpenStudy (anonymous):

Are you remembering to factor the negative across the whole group? Such as -(x+2)=-x-2?

OpenStudy (anonymous):

wait. . . why would that be -x-2?

OpenStudy (anonymous):

Whenever you have a number in front of a group of terms, you must distribute that value across the whole group. For example, \[2(x+1)=2x+2\] Whenever we have a negative, it is the same as using -1, therefore, we have to distribute both the 1 and the negative across the group. \[-1(x+2)=-x-2\]

OpenStudy (anonymous):

In your expression we can simplify the left side as: \[4-(x+2) \rightarrow 4+(-1)(x+2) \rightarrow 4-x-2 \rightarrow 2-x\]

OpenStudy (anonymous):

Thank you! That helped. So the correct answer is x < -7 right?

OpenStudy (anonymous):

X<-7

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

:)